! Suppose that you have a battery with12v, show that a device operating at 4v can be powered by this battery? If estimate the value of resistance that will be used in your proposed cercuit.


1
Expert's answer
2021-12-07T20:11:40-0500

We need a voltage divider:


"R_2=R_1\frac{V_\text{out}}{V_\text{in}-V_\text{out}}=R_1\frac{4}{12-4}=0.5R_1."

For "R_1=10\Omega,R_2=5\Omega."